OpenStudy (anonymous):

What is the equation of a circle with center (3,8) and radius 2?

OpenStudy (zzr0ck3r):

(x-a)^2+(y-b)^2=r^2 (a,b) is the center.

OpenStudy (anonymous):

\[\bf (x-a)^2+(x-b)^2=r^2\]Where \(\bf (a,b)\) is the centre of the circle and \(\bf r\) is the radius. @K_saunders4

OpenStudy (anonymous):

Sorry, it should be \(\bf (y-b)^2\). My apologies.

OpenStudy (anonymous):

\[\bf (x-a)^2+(y-b)^2=r^2\]

OpenStudy (anonymous):

so the answer is (x-3)^2+(y-8)^2=2

OpenStudy (anonymous):

Thanks so much guys

OpenStudy (anonymous):

idk I have options here a-d and none of them says that

OpenStudy (anonymous):

r^2 is 4, not 2

OpenStudy (anonymous):

A. (X-3)^2+(y-8)^2=4, B. (x+3)^2+(y+8)^2=4 or C. (x+3)^2+(y+8)^2=2 or D. (x-3)^2+(y-8)^2=2 which one?

OpenStudy (anonymous):

a) loolololol

OpenStudy (anonymous):

So a?

OpenStudy (anonymous):

Lol sorry and okay thanks so much